# A new indeterminacy-based quantum theory

### v4

@@ 3 Why is EnDQT local?
It doesn’t seem that any current quantum theories, which don’t modify the fundamental equations of QT, can use QCMs in this local, non-relationalist, and non-operational way to give a local common cause explanation of quantum correlations like the ones in the Bell and extended Wigner’s friend scenarios. So, EnDQT seems to be currently the only one to be able to do so in this way. Note that relationalist theories are, along with EnDQT, the only non-operational non-hidden variable theories that don’t modify the fundamental equations of QT and consider it a universal theory. So, they are the only ones who could also consider that QCMs, which use standard QT, provide the whole causal story. Furthermore, spontaneous collapse theories and gravity collapse theories will necessarily impose fundamental limitations on these macroscopic superpositions. As I have said, in the case of EnDQT, it all depends on the details of the histories of the SDCs (which includes not being subject to human ingenuity). However, typically, in relationalist theories, the shared correlations of the friends or Wigners only arise when they meet (if they ever meet). So, there isn’t a common cause explanation in the above sense. Moreover, QCMs in the single-world cases (at least) should be modified or adapted to account for these multiple varying perspectives, since they don’t take into account that variation. So, contrary to the suggestions of others, EnDQT considers that QCMs don’t need to be modified or adapted to a relationalist approach in order for them to explain the correlations that arise in the extended Wigner’s friend scenarios. [^74] Cavalcanti & Wiseman (2021), Schmid et al. (2023), and Yīng et al (2023). See Ormrod & Barrett (2024) for a recently proposed relationalist adaptation. Adopting EnDQT, we don’t need to adopt this view that may be seen as more complex, and this can be regarded as another benefit of this view.
@@
Finally, how to support P-2) and P-3)? QCMs consider that common causes can have indeterminate values represented via QT, i.e., via subsystems of an entangled state, and probabilities explicitly don’t arise from the ignorance of underlying determinate values. Contrary to CCMs and in agreement with EnDQT, for QCMs, common causes, represented by the subsystems of the entangled state, don’t have determinate values and a Born probabilistic model independently of the interactions with Alice or Bob. Also, contrary to CCMs, the relations of influence are explicitly represented via QT, i.e., via CPTP maps when systems don’t interact with members of an SDC and by CP-maps/POVMs when systems interact with members of SDCs. Only in a limit where we can consider the systems as having determinate values, the QMC reduces to the CMC. [^75] The details about how to obtain this limit precisely are too evolved to be presented here. Basically, the classical limit should involve a process operator $\sigma_{A_{1}\ldots A_{n}}$, where there is an orthonormal basis at each node (that is, an orthonormal basis for $\mathcal{H}_{A_{i}^{\text{in }}}$, along with the basis for $\mathcal{H}_{A_{i}^{\text{out }}}$ ), such that $\sigma_{A_{1}\ldots A_{n}}$ is diagonal with respect to the product of these bases. This corresponds, for example, to the situation where the systems at the source are prepared in a product state.
So, according to EnDQT, CCMs with their CMC are inappropriate to provide an account of quantum causal relations, contrary to QCMs. The latter provides an appropriate causal explanation of Bell correlations, which is local. In this way, EnDQT deals with Bell’s theorem.
Sometimes, it is argued that QT is non-local and that the EPR argument (Einstein et al., 1935) ruled out the existence of local indeterministic theories (e.g., Maudlin, 2014), and so one might worry that there is something wrong with my argument above. However, such argument concerning the non-locality of QT shouldn’t be right because EnDQT, as an indeterministic local theory, is a counterexample to that claim. I don’t have space to enter into details, but note that the so-called EPR criterion of reality assumed in this argument can be precisely seen (Gömöri & Hofer-Szabó, 2021) as a consequence of the classical Reichenbach principle, which, as I have mentioned, is a special case of the more general CMC (Hitchcock & Rédei, 2021). However, EnDQT doesn’t consider that the CMC can, in general, represent causal relations between quantum systems. Thus, it rejects the EPR criterion as representing such causal relations and one of the basic premises of that argument for non-locality.
@@ 4 Conclusion and future directions
I have proposed EnDQT and argued that, contrary to the other well-known quantum theories, it has the great benefit of being a local, non-relational, and non-superdeterministic/non-retrocausal QT. Systems have determinate values only while interacting with other systems of SDCs. On top of that, EnDQT has the benefit of being conservative, not modifying the fundamental equations of QT, and, in principle, arbitrary systems can be placed in a superposition for an arbitrary amount of time. Also, EnDQT is able to give a local causal explanation of quantum correlations. There are many future directions. For instance, one should develop more realistic models involving initiators, seek to develop techniques to map SDCs, and test and extract further predictions from EnDQT, which might distinguish it empirically from other quantum theories. [^76] See appendix C for further discussion on future directions.

### v5

@@ 3 Why is EnDQT local?
As a side note, it doesn’t seem that any of the current QTs that don’t modify the fundamental equations of QT, can use QCMs in this local, non-relationalist, and non-operational way to give a local common cause explanation of quantum correlations like those in the Bell and extended Wigner’s friend scenarios. Thus, EnDQT currently appears to be the only theory capable of providing such an explanation in this manner. Note that relationalist theories are, along with EnDQT, the only non-operational, non-hidden variable theories that don’t modify the fundamental equations of QT and consider it a universal theory. So, they are the only ones who could also consider that QCMs, which use standard QT, provide the whole causal story. Furthermore, spontaneous and gravity collapse theories will necessarily impose fundamental limitations on these macroscopic superpositions. As I have said, in the case of EnDQT, it all depends on the details of the histories of the SDCs (which includes not being subject to human ingenuity). However, typically, in relationalist theories, the shared correlations of the friends or Wigners only arise when they meet (if they ever do). Thus, there is no common cause explanation in the above sense. Moreover, QCMs in the single-world cases (at least) should be modified or adapted to account for these multiple varying perspectives since they don’t consider that variation. So, contrary to the suggestions of others, EnDQT considers that QCMs don’t need to be modified or adapted to a relationalist approach for them to explain the correlations that arise in the extended Wigner’s friend scenarios.109109109[24, 90, 106]. See [75] for a recently proposed relationalist adaptation. Adopting EnDQT, we don’t need to adopt this more complex approach to QCMs, which can be regarded as another benefit of this view.
Therefore, I have shown that if QCMs consider that i) systems only assume determinate values with a certain Born probability when they interact with an SDC, and ii) where those relations are described via QT, but without reifying the quantum states like EnDQT assumes, then QCMs provide a local non-relational, non-superdeterministic, and non-retrocausal explanation of quantum correlations.
Finally, let’s turn to my last core premise:
P3) If we interpret Classical Causal Models according to EnDQT, we see that in comparison to QCMs i) they don’t assume that systems only have determinate values when they interact with an SDC and ii) that the relations of influence aren’t described via QT. The relations of influence that they represent only arise in a special limit from QCMs.
QCMs consider that common causes can have indeterminate values represented via QT, i.e., via subsystems of an entangled state, and probabilities explicitly don’t arise from the ignorance of underlying determinate values. Furthermore, in agreement with EnDQT, for QCMs, common causes, represented by the subsystems of the entangled state, don’t have determinate values and a Born probabilistic model independently of the interactions with Alice or Bob. This is contrary to CCMs that assume that common causes and causes in general are represented via classical variables, which can always be assigned probabilities. Therefore, CCMs don’t assume i).
@@
Also, CCMs don’t assume ii) because the relations of influence aren’t explicitly represented via QT, i.e., via CPTP maps when systems don’t interact with members of an SDC and by CP-maps/POVMs when systems interact with members of SDCs. Only in a limit where we can consider the systems as having determinate values, the QMC reduces to the CMC.110110110The details about how to obtain this limit precisely are too evolved to be presented here. Basically, the classical limit should involve a process operator $\sigma_{A_{1}\ldots A_{n}}$, where there is an orthonormal basis at each node (that is, an orthonormal basis for $\mathcal{H}_{A_{i}^{\text{in }}}$, along with the basis for $\mathcal{H}_{A_{i}^{\text{out }}}$ ), such that $\sigma_{A_{1}\ldots A_{n}}$ is diagonal with respect to the product of these bases. This corresponds, for example, to the situation where the systems at the source are prepared in a product state.
Thus, if we interpret classical Causal Models according to EnDQT, we see that i) systems only assume determinate values with a certain Born probability when they interact with members of an SDC and ii) the relations of influence that they represent aren’t described via QT. The relations of influence that they represent only arise in a special limit from QCMs.
Therefore, given that
P4) QCMs, as interpreted by EnDQT, consider that i) systems only assume determinate values with a given Born probability when they interact with an SDC, and ii) where those relations are described via QT, but without reifying the quantum states like EnDQT assumes;
and given P1) and P2), QCMs accurately represent causal relations between quantum systems for EnDQT and where those relations of influence are local. Furthermore, given P1) and P3), contrary to QCMs, CCMs with their CMC for EnDQT don’t accurately represent relations of influence between quantum systems. Contrary to the previous argument, notice that this argument doesn’t necessarily appeal to the derivation by [79] of the CMC to reject the applicability of the CMC to infer causal relations between quantum systems (although it may $also$ appeal to this proof to help to compare both models). Rather, it appeals in general to the interpretation and comparison of CCMs with QCMs according to the EnDQT. Also, it shows that we can have a local explanation of Bell correlation via QCMs, and so it prompts us to reject the CCMs and the CMC.
Coming back to Bell’s theorem, given the above arguments, such as the argument for locality, we reject CCMs and the CMC as accurately representing the relations of influence between quantum systems. Therefore, we reject the Classical Reichenbach Common Cause Principle and, therefore, the factorizability condition as accurately representing the relations of influence between quantum systems. EnDQT rather assumes a local and more accurate way of accounting for those correlations via QCMs. Therefore, via these reasons and arguments, EnDQT deals with Bell’s theorem quantum correlations while ensuring that the theory does not conflict with relativity by favoring any specific reference frame, thus avoiding action-at-a-distance phenomena as seen in Bohmian mechanics. Also, it does not introduce hidden variables that result in retrocausality or superdeterminism. Furthermore, it deals with this theorem without adopting a relationalist interpretation of QT. Therefore, it fulfills desiderata LC) and NR).
One may object that according to EnDQT, systems can also have an indeterminate value of position because systems can be in a superposition of positions or be in an entangled state of positions. Thus, it seems that some sort of non-locality is built into EnDQT because systems won’t have a location in spacetime in the standard sense. As a reply, I should note that at the level of fundamental physics, i.e., at the level of quantum fields and quantum field theory, time and position aren’t observables in the standard sense. They are rather parameters that serve to parametrize the different configurations of quantum fields. So, since non-relativistic quantum theory arises as a limit of relativistic quantum theory (i.e., quantum field theory), we shouldn’t consider position and time per se as observables of quantum systems. So, in particular, we shouldn’t consider a system in a superposition of positions as literally spread out over space at a time. Thus, the above objection is not worrying.
Sometimes, it is argued that QT is non-local and that the EPR argument [37] ruled out the existence of local indeterministic theories (e.g., see [65] for an influential argument), and so one might worry that there is something wrong with my argument above. However, this argument concerning the non-locality of QT in principle should not be correct because EnDQT, as an indeterministic local theory, is a counterexample to its claims.
The EPR argument can be posed in the following way, let’s assume QT and that experiments performed at arbitrary distances from each other, don’t disturb the outcomes of each other. Furthermore, let’s assume the following so-called EPR criterion of reality:111111111[37].
[i]f, without in any way disturbing a system, we can predict with certainty (i.e., with probability equal to unity) the value of a physical quantity, then there exists an element of reality corresponding to that quantity.
Let’s consider Alice and Bob in distantly separated labs, performing measurements in their systems in the same direction. Every time Alice gets spin up, Bob gets spin down, and vice-versa. If Alice/Bob, without disturbing the outcome, can predict with certainty the value of the observable of the system that Bob/Alice measures, then there should exist an element of physical reality that corresponds to these quantities that Bob/Alice measures prior to their measurement. However, QT is silent about those elements for all physical magnitudes in these situations hence it is incomplete.
Now comes another assumption for the above argument for non-locality:
The further conclusion that a final and complete physical theory must be deterministic at least with respect to these particular phenomena just comes as an additional bonus. If the world is EPR-local, and there is no spooky action-at-a-distance, then not only must the quantum mechanical description of a system leave out some elements of reality, but the elements that it leaves out must be sufficient, in these circumstances, to completely predetermine the outcome of the “measurement” operation. For, as Bell remarks in the passage cited above, “any residual undeterminism could only spoil the perfect correlation”. This further conclusion of predetermination obviously requires that the relevant correlations be perfect, which is also what is required here to apply the EPR criterion (“we can predict with certainty (i.e. with probability equal to unity)”). ([65], pp. 10)
Thus, this completion should be a deterministic completion. Bell’s 1964 theorem112112112[9]. formalizes this conclusion by an assumption, which we may call outcome determinism or counterfactual definiteness or predeterminism. Let’s consider that $c$ is the variable that designates the preparation at the source of entangled particles, $\lambda$ represents the hidden variables of the particles that may exist between the source and the measurement of Alice and Bob. $A$ and $B$ are the measurement outcomes of Alice and Bob, and $a$ and $b$ are their choices of measurement settings. $P_{\theta}$ designates a probability model given by a theory $\theta$. Then, a theory $\theta$ obeys predeterminism if assumes variables $\lambda$ that determine the outcome A and B in such a way that:113113113See [102].
$\forall A,a,B,b,\lambda,\ P_{\theta}(A,B\mid a,b,c,\lambda)\in\{0,1\}.$ (15)
We consider that these variables fully determine Alice and Bob’s outcomes. In addition to this assumption and a no-superdeterminism assumption, this theorem also assumes the locality condition, also known as “parameter independence”114114114[52], which says that the outcome for the quantum system in one of the wings in the Bell scenario is statistically independent of the measurement performed in the other wing and vice-versa, i.e., $P_{\theta}(B\mid a,b,c,\lambda)=P_{\theta}(B\mid b,c,\lambda)$ and similarly in the other case. The argument for non-locality proceeds by claiming that given that we should accept the EPR argument (as Bell supposedly did), we should already accept the predeterminism assumption. 115115115As Maudlin (2014, pp. 10) argues, That argument is one line: the very “element of reality” that the EPR argument proves to exist—given EPR-locality—is an element of reality defined just as whatever physical characteristic of the system it is that ensures how it would react to the measurement in question. So any system that has that element of reality has a physical characteristic that determines how it would react to the measurement. But that just is determinism with respect to that particular “measurement operation”. And the EPR argument can be repeated for any “measurements” for which quantum theory predicts perfect correlations between the outcomes and that can be made arbitrarily far apart in space. Hence, in an EPR- local theory both the reactions to a “position measurement” and to a “momentum measurement” must be predetermined by some element of reality in the system, and in the Bohm spin example the reactions to every possible “spin measurement” must be predetermined. That is enough to get Bell’s 1964 argument off the ground. Not by assuming determinism, but by assuming EPR-locality and deriving determinism.
Thus, since denying the “no-superdeterminism” and the “single outcomes” assumption also assumed by this no-go theorem is problematic,116116116[65]. we should rather reject the locality assumption. Hence, quantum theory is non-local, i.e., it leads to action at a distance. This argument for non-locality is summarized in Figure 6.
Figure 6: Diagram that helps understand the argument for non-locality based on Bell’s 1964 theorem works and how EnDQT deals with it. It again involves the rejection of the applicability of the Causal Markov Condition and Classical Causal Models to, in general, accurately represent causal relations between quantum systems. If a box with the claim $X$ is connected by one or more arrows without a legend, means that the one or more boxes with claims $Y$ that point to $X$ jointly lead to the derivation of $X$. If an arrow has the legend “lead to,” it means that one claim leads to another, as explained in the text.
There are some problems with this argument.
First, Bell’s 1964 theorem is restricted to perfect correlations that occur in certain circumstances in the Bell scenario when Alice and Bob measure their entangled systems on the same basis. Predeterminism (equation (15)) assumed by this theorem is that there is a theory that allows us to predict with certainty the outcomes of Alice and Bob by specifying certain variables in the past of these outcomes plus the measurement choices of Alice and Bob. These variables are any information that represents entities that determine that outcome.
The EPR criterion of reality is about the postulation of an inference from certain predictions to something that determines those predictions. Predeterminism isn’t that assumption, and an indeterministic theory like EnDQT doesn’t obey it.117117117It is also unclear if Alice can ever predict with $100\%$ of probability the outcome of Bob and vice-versa in a world represented by an indeterministic theory. The particles that constitute Bob obey indeterministic dynamics, and thus, it is impossible to be certain whether Bob will measure his particle on the same basis as Alice, as it is assumed by the EPR criterion. So, one could argue that the EPR criterion assumes physically unrealistic circumstances for an indeterministic theory.
Second, it is not even clear to what extent this assumption follows from the EPR criterion of reality (CR) since it is unclear what this principle means. To my knowledge, the best precisification of this principle views it as a consequence of the classical Reichenbach common cause principle [45], which, as I have mentioned, is a special case of the more general CMC applied in realistic physical situations [49].
First of all, instead of variables, [45] express the CRCCP using events described using propositions and their negation. However, in the case of their proof, in principle, we can think of these propositions in terms of binary variables. Now, some terminology: there exists a directed path from vertex $X$ to vertex $Y$, denoted by $X\rightsquigarrow Y$, if there is a direct arrow from $X$ to $Y$ ($X\rightarrow Y$) or there is an intermediate node $Z$ such that $X\rightsquigarrow Z$ and $Z\rightarrow Y$. Furthermore, we write $X\nrightarrow Y$ when event $X$ doesn’t influence event $Y$.
Suppose A is the outcome of a measurement $a$ of a certain physical quantity measured by Alice. Suppose that B denotes the event that involves the prediction that the quantity in question has a certain value, which is made by Bob. Furthermore, suppose event B doesn’t influence event A, as it should, because the act of prediction doesn’t influence what is being predicted. Also, suppose that event B doesn’t influence event A. In other words, suppose that Bob doesn’t influence the outcome of Alice, i.e.,“without in any way disturbing a system” the outcome of Alice. Also, suppose that the prediction involved in event B is certain, and by certain, it’s meant that whenever B happens and a is performed, the measurement results in outcome A, i.e., $p(A|a\land B)=1$. Then, there is an element of reality $\alpha$ that will cause the outcome A to occur when the measurement of a is performed. In other words, $p(A|a\land\alpha)=1$. Notice that the notion of determination of an outcome by elements of physical reality is clearer here. We can summarise the above description in the following way ([45], pp. 13459-13460):
Reality Criterion Suppose that $A$ is an outcome of measurement $a$. Assume that there is an event $B$ such that
$A\nrightarrow B\text{ and }B\nrightarrow A$ (24)
$p(A|a\land B)=1$ (25)
Then there exists a further event $\alpha$ such that
$\alpha\rightarrow A$ (26)
$p(A|a\land\alpha)=1$ (27)
[…]
Common Cause Principle Suppose that $A$ is an outcome of measurement $a$. Assume that there is an event $B$ such that
$A\nrightarrow B\text{ and }B\nrightarrow A$ (33)
$p(A\land B|a)\neq p(A|a)p(B|a)$ (34)
Then, there exists a further event $C$ such that
$C,\neg C\rightarrow A,B$ (35)
$p(A\land B|a\land C)=p(A|a\land C)p(B|a\land C)$ (36)
$p(A\land B|a\land\neg C)=p(A|a\land\neg C)p(B|a\land\neg C)$ (37)
$C$ in the equation above would be the element of reality. If we take into account this definition of the CRCCP and the above one of the RC, it can be shown that the reality criterion is just a special case of the CRCCP. More concretely, it can be shown that the reality criterion is an application of the CRCCP to the case where the prediction is certain and if the act of prediction does not disturb the predicted phenomenon.
So, even if we grant the above connection between the EPR criterion of reality and predeterminism since EnDQT rejects the CRCCP’s applicability to accurately represent causal relations between quantum systems, it also rejects the CR’s applicability to represent such causal relations (since one follows from the other). Thus, (as expected) it should also reject the predeterminism condition as accurately representing such relations.
Maudlin120120120[65]. called the reality criterion analytic, but as Lewis puts it
I doubt that Einstein’s criterion of reality is really analytic […] It seems perfectly conceivable that an event could be predicted with certainty even when there is nothing physical that brings that event about. […] Indeed, Maudlin is perfectly sanguine about fundamentally probabilistic laws (e.g. in spontaneous collapse theories), according to which there is in general no physical reason why this result is obtained (as opposed to that result) when the probabilities differ from zero and one. Why should things be different when the probabilities are zero and one? ([59], pp. 38–39)
Gömöri and Hofer-Szabó reply:
[t]he answer to this is now simple and clean: the reason why probability-1 predictions are distinguished is because those predictions mean perfect correlation between the act of prediction and the predicted outcome, and a perfect correlation can only be explained, in accord with the CCP’s requirement [i.e., the CRCCP], by a deterministic common cause; hence the outcome must be predetermined. ([45], pp. 13463).
If the CRCCP and the CMC, more generally, are based on determinism, and this seems to be the case given what we saw above with one justification of the CMC, this is unsurprising. Moreover, as my arguments above indicate, quantum causal models show how the CRCCP leads to wrong inferences about how systems influence each other. They also show us how Bell’s theorems, more generally, can be considered as telling us how quantum indeterminacy and the determinacy that arises from it behave and that we need new inferential tools to understand this process. More concretely, the inferences we make based on classical causal models (and their assumptions) are wrong for EnDQT, even when perfect correlations arise between measured quantum systems. When Alice, at one wing of a Bell experiment, makes her measurement, she can’t make claims about the physical state of Bob’s system a la elements of reality/hidden variables, which will determine his result. She is applying the wrong causal model to make these inferences. The right theories to use to make inferences are quantum causal models and EnDQT in this view. If Alice used them, she would see that not only her target system but also Bob is an indeterministic system connected with SDCs that determine his behavior. Furthermore, Bob, as well as his target system, are described via quantum states, observables, as well as SDCs and QCMs. She would also notice that there is no “element of physical reality” with the above features in the quantum domain, and nothing is predetermined in Bell’s theorem sense.
Thus, for EnDQT, the above argument doesn’t work to show that QT is non-local, even if we grant that we can go from this argument to accepting predeterminism (which we can also deny on independent grounds, as well as the predeterminism assumption, as I have explained above).
One may object to the claim that EnDQT provides a local view of the world and argue that there seems to be some non-locality going on in this view. Suppose that quantum systems have some kind of irreducible and intrinsic dispositions that are manifested with a given probability under some measurement-like interactions. Also, quantum states work as predictors for those manifestations. Let’s consider the case where the two entangled spin-1/2 particles end up being anti-correlated in the Bell scenario upon measurements of Alice and Bob on the same basis. For instance, it seems that when Bob measures his system and obtains spin-up, the probabilistic disposition of the particle of Alice changes instantaneously at a distance. More concretely, it changes to $100\%$ being spin-down under a measurement of Alice on the same basis as Bob. However, note that it doesn’t seem that any of the theorems or arguments we have examined and assumptions that we have dealt with oblige us to infer that there is some non-local influence between the dispositions of both particles. Furthermore, the view taken here on these theorems and arguments doesn’t lead to the commitment of some sort of irreducible and intrinsic dispositions.
Relatedly, notice that,121121121Thanks to Peter Lewis for pressing me to clarify this. even if we have theories where like EnDQT systems don’t have determinate values (described by hidden variables) until they reach Alice or Bob, EnDQT via its theoretical posits and tools (such as SDCs and QCMs) shows that we don’t need to pose those influences and adopt these theories. Let’s consider the Ghirardi–Rimini–Weber (GRW) theory with flashes (determinate values of position) that arise from the collapse of the wavefunction randomly in spacetime. Flashy GRW doesn’t (at least necessarily) adopt a realist position about the wave function – the wave function can be regarded as a bookkeeping device that allows us to keep track of the probabilities of flashes at various locations (i.e., of dispositions). Suppose Alice in a Bell scenario measures the spin of her particle. That triggers a series of flashes at Alice’s location, and those flashes instantly and non-locally change the probabilities for flashes associated with Bob’s particle. For example, if Alice measures z-spin and gets spin-up, the probabilities at Bob’s location instantly change to those associated with a z-spin-down particle. But flashy GRW doesn’t postulate any hidden variables (in the sense of determinate values) underlying the probabilities – they can appeal just to brute dispositions to account for these correlations.
However, note that EnDQT provides a theory that shows how we can also dispense with posing these non-local dispositions while explaining Bell correlations. According to EnDQT, systems represented via QCMs evolve locally and only have determinate values when they interact with members of SDCs. So, SDCs are necessarily involved in these influences that give rise to determinate values, and they concern local interactions between systems, not non-local ones. Alice doesn’t influence non-locally the probabilities of the system of Bob or vice-versa. Thus, EnDQT provides a local $causal$ pathway that dispenses with the adoption of non-local dispositions.
@@ 4 Conclusion and future directions
I have proposed EnDQT and argued that contrary to the other well-known QTs, it has the great benefit of being a local, non-relational, non-superdeterministic, and non-retrocausal quantum theory. Systems have determinate values only while interacting with other systems of SDCs. On top of that, EnDQT has the benefit of being conservative, not modifying the fundamental equations of QT, and, in principle, arbitrary systems can be placed in a coherent superposition for an arbitrary amount of time. Also, EnDQT is able to give a local causal explanation of quantum correlations.
There are many future directions. For instance, one should develop realistic quantum field theoretic models of SDCs,122122122[82]. explore the explanatory potential further, seek to develop techniques to model and map SDCs, and test and extract further predictions from EnDQT, which might distinguish it empirically from other QTs.